mxddi3:

how do you factor 3x^2 -7x + 2 (algebra 1)

jhonyy9:

do you can solve this quadratic equation ?

mxddi3:

im just confused bc the teacher said to do a*c which in this case is 6 and when you factor out positive 6, idk how you make that equal -7 

jhonyy9:

can you post here your work please ?

mxddi3:

yes one second

jhonyy9:

3x^2 = 2x^2 +x^2 -7x = -6x -x 2 = 1 +1 just like an idea

jhonyy9:

ok i wait please post it

mxddi3:

3x^2 - 7x +2 A-term: 3x^2 B-term:-7x C-term: +2 Split term into 2: a*c 3*2=6 Factor out 6: 1*6, 2*3, -1*-6, -2*-3 Must equal the b-term- -7 -1+-6=-7 Now, 3x^2 -x -6x +2 (3x^2-x)(-6x+2) Find what they have in common:

mxddi3:

and this is where im confused

jhonyy9:

Must equal the b-term- -7 -1+-6=-7 -1 +6 not equal -7

mxddi3:

its supposed to say -6

jhonyy9:

ohh yes i see it now

jhonyy9:

(3x^2-x)(-6x+2) Find what they have in common: after this so this is so easy 3x^2 -x what are common factors ?

mxddi3:

x?

jhonyy9:

yes sure this is easy - courage please i m here to help you so that you factoriz out the x what you get ?

mxddi3:

so you would have x and what's left is -1,3,and another x so it would be x(x-3) ?

jhonyy9:

no than you multiplie back the x(x-3) = ?

mxddi3:

she didn't tell us to do that, she said to pull out what they have in common, in this case "x" and then in () put what's left, then do the same thing for the other group and she said the final answer is written like this (this is from one we did yesterday): (y+15)(y-2) so i'm not quite sure what you're asking here, sorry.

jhonyy9:

yes sure i understand but what you made above not is correct factorized out x(x-3) not equal 3x^2 -x

mxddi3:

so then how do you make it correct?

jhonyy9:

how you factoriz out the x from 3x^2 -x ? 3x^2 = 3x*x -x = -1*x yes ? so when you factoriz out an x from these terms what will remain there ?

mxddi3:

im confused....

jhonyy9:

no this is easy 3x^2 -x = 3x*x -1*x = x(?) from these terms of 3x*x and -1*x than you factoriz out the x what will remain there ?

jhonyy9:

dont understand ? please

mxddi3:

your wording is confusing me, sorry. if i factor the x, idk what it does to the equation. i'm still confused

jhonyy9:

i will try explain it to you 3x^2 -x = 3x*x -1*x = x(?) looke please from 3x*x than you factoriz out an x remain there 3x yes ? and from -1*x when you factoriz out an x remain there an -1 yes so all together 3x^2 -x = x(what ? )

mxddi3:

uhm, would it be -3x ? because in the other ones you took away the one x right? and then you had 3x and -1 left so i'm assuming you put those together? i have no idea

jhonyy9:

yes so 3x^2 -x = x(3x-1) ok ?

jhonyy9:

and now the second one (3x^2-x)(-6x+2)

mxddi3:

what do i do here?

jhonyy9:

you need factoriz out the common factor

jhonyy9:

-6x +2 = what is common factor there ?

mxddi3:

2

jhonyy9:

yes so factoriz it out please

mxddi3:

so, we would have 2 and left over is x-3 so i have 2(x-3) ?

jhonyy9:

2(x-3) = 2x -6 this is what was there above ?

jhonyy9:

there is -6x +2 yes ?

mxddi3:

yeah i think

jhonyy9:

so and now how will be correct ?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 -6x = -2*3x and 2 = -2*(-1) what is common factor there ? \(\color{#0cbb34}{\text{End of Quote}}\)

mxddi3:

-2?

jhonyy9:

yes sure from these -2*3x and -2*(-1) you can factorize out the -2 and what will get ?

jhonyy9:

-6x+2 = -2*3x -2(-1) = -2(?)

jhonyy9:

from -2*3x than you factorize out the -2 what remain ?

jhonyy9:

and so too than you factorize out the -2 from -2(-1) what remain ?

mxddi3:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 from -2*3x than you factorize out the -2 what remain ? \(\color{#0cbb34}{\text{End of Quote}}\) 3x?

mxddi3:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 and so too than you factorize out the -2 from -2(-1) what remain ? \(\color{#0cbb34}{\text{End of Quote}}\) -1? or do i multiply them

jhonyy9:

yes

jhonyy9:

and now all together -6x+2 = -2*3x -2(-1) = -2(?)

mxddi3:

well, if i take out the -2's, you have 3x(-1) so does that mean -3x?

jhonyy9:

no because there is between -3x and -1 an minus not multiplication so will be -2(3x-1)

mxddi3:

oh ok. so it's 3x-1 then, is that what you're saying?

jhonyy9:

yes - so now we go back there up and you need writing all together what we have got

jhonyy9:

do you remember on ?

mxddi3:

i had to go reread it, sorry. u, so i got (3x-1)(x-2) is that right?

mxddi3:

uh* not u

jhonyy9:

yes this is right but do you understand now the way ?

jhonyy9:

hope was helpfully have a nice day bye bye